# Dividing Indivisible Items for the Benefit of All: It is Hard to Be Fair Without Social Awareness

### v1

@@ 3 Socially Unaware Agents
For any $\mathcal{F}\in\{\text{tEF1},\text{EF1},\text{sEF1},\text{wEF1},\text{swEF1},\\ \text{EFL},\text{EF}\}$, an instance $\mathcal{I}=({{N}},{M},({v}_{i})_{i\in{{N}}},(s_{i})_{i\in{{N}}})$ admits a SIM and $\mathcal{F}$-fair allocation if and only if an instance $\mathcal{J}=({{N}},{M},({v}_{i})_{i\in{{N}}},({s}^{*}_{i})_{i\in{{N}}})$, where ${s}^{*}_{i}(g)=1$} if $i\in\arg\max_{i\in{{N}}}{s}_{i}(g)$} and 0 otherwise, admits a SIM and $\mathcal{F}$-fair allocation.
With the previous lemma in hand, we show our main hardness result of this section, which settles that even if both the valuations and the social impact functions are binary, the problem is NP-complete. Our strategy in the reduction is to introduce some initial unfairness between agents so that the rest of the items have to be allocated in an envy-free way, which is known to be computationally hard (Aziz et al. 2015; Hosseini et al. 2020).
@@
For   any $\mathcal{F}\,\,\in\,\,\{\,\,\text{tEF1},\,\,\text{EF1},\,\,\text{sEF1},\,\,\text{wEF1},\\ \text{swEF1},\text{EFL}\}$, it is NP-complete to decide whether a SIM and $\mathcal{F}$-fair allocation exist, even if both the valuations and the social impact functions are binary.
Note that in the previous results, the social impact functions and the valuations are not the same. If we additionally assume that the valuations and social impacts are the same, we obtain a positive result modifying the Yankee-swap algorithm of Viswanathan and Zick (2025) and the weighted round-robin of Chakraborty et al. (2021), respectively.
@@
For   any $\mathcal{F}\,\,\in\,\,\{\,\,\text{tEF1},\,\,\text{EF1},\,\,\text{sEF1},\,\,\text{wEF1},\\ \text{swEF1},\text{EFL}\}$, it is NP-complete to decide whether a SIM and $\mathcal{F}$-fair allocation exist, even if the social impact functions are binary and n=2{{n}}=2.
@@
However, the previous hardness result is highly dependent on the fact that the agents’ valuations for the items are exponential in the input size. That is, the result shows only weak NP-hardness. In the following, we complement the hardness and show that if the number of agents is a fixed constant and the valuations are polynomially bounded in the input size, then there is an efficient algorithm. Notably, the algorithm is not based on an explicit DP formulation, as one might expect, but rather models our problem as graph reachability. Thanks to this, the algorithm can be implemented very efficiently by generating the graph on-demand and finding the solution using, e.g., a variant of the A algorithm (Hart, Nilsson, and Raphael 1968; Dechter and Pearl 1985; Korf 1997; Zhou and Zeng 2015).
@@
4 Socially Aware Agents
@@
Motivated by the experimental work of Herreiner and Puppe (2009) and Hosseini et al. (2025), which explore what real-life users perceive to be fair outcomes, Hosseini (2024) argues that “developing fair algorithms may require axioms that are able to capture solutions that take the social context into account beyond perceived envy.” In this section, we build up on this idea and follow the approach introduced by Flammini, Greco, and Varricchio (2025) and assume socially aware agents. Intuitively, an agent is socially aware if they are willing to accept a formally unfair outcome in case it leads to a greater social impact.
@@
Flammini, Greco, and Varricchio (2025) proved that there is a modification of the famous envy-cycle elimination algorithm (Lipton et al. 2004) that always finds an SA-EF1 allocation. We extend their result in two ways. In our first result, we show that SA-swEF1 allocations are also guaranteed to exist. Interestingly, our approach is based on a picking sequence similar to the round-robin algorithm. More precisely, the algorithm is a modification of the weighted picking sequence protocol proposed by Chakraborty et al. (2021), which reduces to a round-robin approach in the case of identical agent weights. However, we only allow agents to pick goods for which they maximize social impact, and we force agents to skip their turn whenever they do not maximize the social impact of any unallocated good. The proof then basically follows from Chakraborty et al. (2021) and a simple observation that agent $i$ can SA-envy agent $j$ only if $j$ chose only goods for which $i$ maximizes social impact. Therefore, from the point of view of $i$, we can completely ignore all agents that get an item for which $i$ does not maximize social impact. On the other hand, if we consider an auxiliary instance in which agents only value items for which they maximize impact and get a dummy item that everyone values 0 whenever they skip a turn we get an allocation that is swEF1 according to Chakraborty et al. (2021) and in which: 1) All agents get same bundle plus some dummy items. 2) Every agent values every bundle for which the agent maximizes the social impact and all non-dummy items in it exactly the same.

### v2

@@ 3. Socially Unaware Agents
For any $\mathcal{F}\in\{\text{tEF1},\text{EF1},\text{sEF1},\text{wEF1},\text{swEF1},\text{EFL},\text{EF}\}$, an instance $\mathcal{I}=({{N}},{M},({v}_{i})_{i\in{{N}}},(s_{i})_{i\in{{N}}})$ of fair division with social impact admits a SIM and $\mathcal{F}$-fair allocation if and only if an instance $\mathcal{J}=({{N}},{M},({v}_{i})_{i\in{{N}}},({s}^{*}_{i})_{i\in{{N}}})$, where ${s}^{*}_{i}(g)=1$} if $i\in\arg\max_{i\in{{N}}}{s}_{i}(g)$} and 0 otherwise, admits a SIM and $\mathcal{F}$-fair allocation.
Proof.
For our problem, we have to maximize the sum of social impact in a solution; regardless of the fairness notion/solution concept. Hence, we can consider any $\mathcal{F}$ of $\mathcal{F}$-Fair Division with Social Impact for this (since the part of the instance that is changing is the social impact “scores”).
Consider that we have a solution for $\mathcal{I}$. In a SIM solution, agents who maximize the social impact for a single item must receive this in the allocation because the social impact function is additive. In the same allocation in the corresponding instance $\mathcal{J}$, the solution must also be SIM because some agent $i$ can only receive some item $g$ if they are a social impact maximizer, i.e., $s^{*}_{i}(g)=1$. Which, by definition, means they are the social impact maximizer in $\mathcal{I}$.
On the other hand, consider we have a solution for instance $\mathcal{J}$ where the social impact function is binary. Consider now that finding a SIM solution in the binary instance did not give us a SIM assignment in the original instance. Therefore, this means there exists some agent $i$ that has a higher social impact for an item $g$ than agent $j$ who received it. If this was the case, agent i would not have been assigned $v_{i}(s)=1$ for $g$ in the binary instance and hence could not have received the item. Hence, contradiction. Therefore it must be a solution in $\mathcal{I}$.
Note that, in both directions, the valuation function $v_{i}$ of agent $i$ remains the same. Therefore, if we have a solution in $\mathcal{I}$ which is $\mathcal{F}$-fair the same allocation is also $\mathcal{F}$-fair in $\mathcal{J}$, vice versa, because the fairness notion only depends on the agents’ valuations. ∎
With the previous lemma in hand, we show our main hardness result of this section, which settles that even if both the valuations and the social impact functions are binary, the problem is NP-complete. Our strategy in the reduction is to introduce some initial unfairness between agents so that the rest of the items have to be allocated in an envy-free way, which is known to be computationally hard (AzizGMW2015, 8, 37).
@@
For any $\mathcal{F}\in\{\text{tEF1},\text{EF1},\text{sEF1},\text{wEF1},\text{swEF1},\text{EFL}\}$, it is NP-complete to decide whether a SIM and $\mathcal{F}$-fair allocation exist, even if both the valuations and the social impact functions are binary.
Proof.
(AzizGMW2015, 8, Theorem 11 ), show that it is NP-hard to find an envy-free allocation of items even when agents have binary additive valuations.
In an instance of envy-free allocation $\mathcal{I}$ we have a set of items $M$ and a set of agents $N$. The goal is to find an envy-free allocation.
From $\mathcal{I}$ we create an instance $\mathcal{J}$ of $\mathcal{F}$-Fair Division with Social Impact. We create some additional items to add to the set $M$. We will call the set of these special items $M^{\prime}$. For each agent $i\in N$ we create an item $g_{i}$ with $v_{i}(g_{i})=0$ and $s_{i}(g_{i})=1$. For all other agents $i\neq j$, let $v_{j}(g_{i})=1$ and $s_{j}(g_{i})=0$.
As for the remaining “standard” items, i.e., the set $M$, the valuation functions remain the same. For each agent $i\in N$ let $s_{i}(g)=1$ for all $g\in M$.
Observe that in the solution, every agent $j$ must receive his special items $g_{j}$ because he is the only agent who maximizes social impact for these. Note that in a SIM solution, any agents can receive the items in $M$. as they all have social impact 1.
EF1   Given a solution to $\mathcal{I}$, we have an allocation of the items in $M$ which is envy-free; what remains to show is that by assigning items the special items in $M^{\prime}$ we still end up with an EF1 allocation. Since we must have a solution which is SIM and EF1, all special items must be given to their respective agents for it to be SIM, i.e., we give item $g_{i}$ to the agent $i$ who is the only agent with $s_{i}(g_{i})=1$. This is clearly an EF1 allocation, as no agent was envious prior to this assignment and every agent has received exactly one item. (Therefore, if any agent became envious, it must be envy-free after removing one good.)
On the other hand, take a solution to $\mathcal{J}$. Each special item must be allocated to its respective agents, since we are given a SIM solution. Without loss of generality, we can assume that $M$ is non-empty. For each agent $i$, he must get allocated at least one standard item in addition to $g_{i}$ or else he will EF1-envy (i.e., EF1 is not satisfied for him) as his bundle will be of value 0 for him. Thus, this allocation of items in $M$ must be envy-free, because already due to the special items agent $i$ has to remove $g_{j}$ in order not to envy $j$, i.e., they cannot afford to envy agent $j$ for any more items.
EFL, sEF1,wEF1,swEF1,   Since we have binary valuations, EF1 and EFL coincide. Strong EF1 requires all agents who envy agent $j$ to be removing the same item. The above construction straightforwardly holds for sEF1, as $g_{i}$ can always be the item that agents are removing. In addition, EF1 is a special case of wEF1 (likewise, sEF1 for swEF1); therefore, the construction follows.
tEF1   We make a small modification to the construction. For each agent, we create two copies of the special item, referred to $g_{i}^{1}$ and $g_{i}^{2}$. The key difference is that there are now two items that must go to each agent, and every other agent will envy them.
Let $v_{i}^{o}(A_{i})$ be agent $i$’s valuation function for instance $\mathcal{I}$ where he receives the bundle $A_{i}$. Let $v_{i}^{r}(A_{i}^{\prime})$ be agent $i$’s valuation function in instance $\mathcal{J}$ where he receives the bundle $A_{i}^{\prime}$.
In our construction, we have: $v_{i}^{o}(A_{i})=v_{i}^{r}(A_{i}^{\prime})$ and $v_{i}^{o}(A_{j})=v_{i}^{r}(A_{j}^{\prime})-2$ as in $\mathcal{I}$ they have received $g_{i}^{1}$ and $g_{i}^{2}$.
Moreover, given a solution to $\mathcal{I}$, we have that $v_{i}^{o}(A_{i})\geq v_{i}^{o}(A_{j})$. Therefore by giving out the special items to their respective agents, we get a SIM solution which is tEF1 because the following holds: $v_{i}^{r}(A_{i}^{\prime}\cup\{g_{i}^{1}\}=v_{i}^{o}(A_{i})+1\geq v_{i}^{o}(A_{j})+1=v_{i}^{r}(A_{j}^{\prime}\setminus\{g_{j}^{1}\})=v_{i}^{r}(A_{j}^{\prime})-1$.
Given a solution to $\mathcal{J}$, we have a SIM solution where ${g_{i}^{1},g_{i}^{2}}\in A_{i}^{\prime}$. Hence, for a pair of agents $i$ and $j$ the following holds because the allocation is tEF1:  $v_{i}^{o}(A_{i})=v_{i}^{r}(A_{i}^{\prime}\cup\{g_{j}^{1}\}-1)\geq v_{i}^{r}(A_{j}^{\prime}\setminus\{g_{j}^{1}\})-1=v_{i}^{o}(A_{j})+1-1=v_{i}^{o}(A_{j})$~, i.e., agent $i$’s valuation for $A_{i}^{\prime}$ is greater than his valuation for $A_{j}$ if he transfers $g_{j}^{2}$, therefore $v_{i}^{o}(A_{i})\geq v_{i}^{o}(A_{j})$ (envy-free). ∎
Note that in the previous results, the social impact functions and the valuations are not the same. If we additionally assume that the valuations and social impacts are the same, we obtain a positive result modifying the Yankee-swap algorithm of (viswanathan2023yankee, 49) and the weighted round-robin of (ChakrabortyISZ2021, 21), respectively.
@@
Proof.
EF1, EFL, sEF1, tEF1.   In their paper, (viswanathan2023yankee, 49) prove that an EFX [^2] For this result, we stress here that (viswanathan2023yankee, 49) assume EFX definition where envy can be eliminated by the removal of any item (even those valued 0 by the agent) as defined by (PlautR2020, 44). This notion is sometimes also referred to as EFX0. allocation with maximum utilitarian social welfare can be found in polynomial time using the algorithm they introduce: Yankee Swap. Observe that it is SIM because $v_{i}=s_{i}$, i.e., a maximum allocation with respect to the agents’ valuation functions is also a SIM solution. Thus, for our problem, because EFX $\implies$ EFL $\implies$ EF1 $\implies$ tEF1 and EFX $\implies$ sEF1, we have existence and a polynomial time algorithm.
wEF1, swEF1.   In their paper, (ChakrabortyISZ2021, 21) introduce an algorithm to find swEF1 allocations in polynomial time. In this algorithm, agents select items in a weight-based picking sequence, and after each round, the partial allocation is swEF1. No agent ever receives an item of value 0, as they can stop picking once there is no item that they value. Hence, since $v_{i}=s_{i}$ the allocation is SIM. Since swEF1 $\implies$ wEF1, we have existence and a polynomial time algorithm for both notions. ∎
@@
For any $\mathcal{F}\in\{\text{tEF1},\text{EF1},\text{sEF1},\text{wEF1},\text{swEF1},\text{EFL}\}$, it is NP-complete to decide whether a SIM and $\mathcal{F}$-fair allocation exist, even if the social impact functions are binary and n=2{{n}}=2.
Proof.
Our reduction is from the Partition problem. There, we are given a multi-set of integers $W=\{w_{1},\ldots,w_{\ell}\}$ such that $\sum_{j\in[\ell]}w_{j}=2t$ and the goal is to decide whether a set $J\subseteq[\ell]$ exists so that $\sum_{j\in J}w_{j}=\sum_{j\in[\ell]\setminus J}w_{j}=t$. The problem is known to be NP-hard, even if no $w_{j}$$j\in[\ell]$, is greater than or equal to $t$ (GareyJ1979, 31).
Given an instance $\mathcal{J}$ of Partition, we construct an equivalent instance $\mathcal{I}$ of of fair division with social impact as follows (for the sake of exposition, we start with the case of EF1 allocations; we show how the reduction works for other fairness notions assumed at the end of the proof). There are two agents, $a_{1}$ and $a_{2}$. The set of items M{M} contains one item $g_{j}$ for every integer $w_{j}\in W$ and two additional large items $G_{1}$ and $G_{2}$. The valuations and social impacts are as follows. For any $a_{i}$$i\in[2]$, we have ${v}_{a_{i}}(g_{j})=w_{j}$ and ${s}_{a_{i}}(g_{j})=1$. Additionally, we have ${v}_{a_{1}}(G_{1})={v}_{a_{2}}(G_{2})=0$${s}_{a_{1}}(G_{1})={s}_{a_{2}}(G_{2})=1$, and ${v}_{a_{1}}(G_{2})={v}_{a_{2}}(G_{1})=t$ and ${s}_{a_{1}}(G_{2})={s}_{a_{2}}(G_{1})=0$. The valuation functions and social impact functions are also summarized in the following table.
 $G_{1}$  $G_{2}$  $g_{j}$$j\in[\ell]$
 $a_{1}$  s=1,v=0{s}=1,{v}=0  s=0,v=ts=0,{v}=t  $s=1,{v}=w_{j}$
 $a_{2}$  s=0,v=t{s}=0,{v}=t  s=1,v=0s=1,{v}=0  $s=1,{v}=w_{j}$
The idea again is that the large items introduce some initial envy, so the rest of the items have to be split equitably (or, in an envy-free way, to use the terms of our problems) between the agents.
For correctness, let $J$ be a solution for $\mathcal{J}$. We construct an allocation ${\mathcal{A}}=({A}_{1},{A}_{2})$ such that ${A}_{1}=\{G_{1}\}\cup\{g_{j}\mid j\in J\}$ and ${A}_{2}=\{G_{2}\}\cup\{g_{j}\mid j\in[\ell]\setminus J\}$, and claim that $\mathcal{A}$} is EF1 allocation. For agent $a_{1}$, we have ${v}_{a_{1}}({A}_{1})=t$ and ${v}_{a_{1}}({A}_{2})=2t$; however, the envy can be eliminated by the removal of $G_{2}$. The argument for $a_{2}$ is symmetric. That is, $\mathcal{A}$} is EF1 and is obviously SIM.
In the opposite direction, let ${\mathcal{A}}=({A}_{1},{A}_{2})$ be an EF1 and SIM allocation. Since $\mathcal{A}$} is SIM, it must be the case that $G_{1}\in{A}_{1}$ and $G_{2}\in{A}_{2}$. Now, we construct a solution for $\mathcal{J}$ by setting $J=\{j\in[\ell]\mid g_{j}\in{A}_{1}\}$. It remains to show that $J$ is indeed a solution. For the sake of contradiction, let $\sum_{j\in J}w_{j}<\sum_{j\in[\ell]\setminus J}w_{j}$ and $\delta$ be the difference between these two sums. Then, in $\mathcal{I}$, it holds that ${v}_{a_{1}}(A_{2})=t+t+\delta=2t+\delta$, while ${v}_{a_{1}}(A_{1})=t-\delta$. Even if we remove the best item from ${A}_{2}$, the envy from $a_{1}$ towards $a_{2}$ persists. This contradicts that $\mathcal{A}$} is an EF1 allocation. Hence, $\sum_{j\in J}w_{j}\geq\sum_{j\in[\ell]\setminus J}w_{j}$. If $\sum_{j\in J}w_{j}>\sum_{j\in[\ell]\setminus J}w_{j}$, then by symmetric arguments we obtain that $a_{2}$ has EF1-envy towards $a_{1}$. Thus, it must be the case that $\sum_{j\in J}w_{j}=\sum_{j\in[\ell]\setminus J}w_{j}$, which shows that $J$ is indeed a solution for $\mathcal{J}$.
sEF1, wEF1, swEF1.  In case of two agents, sEF1 and EF1 coincides. Moreover, EF1 is a special case of wEF1 where all weights are the same, and the same is true for sEF1 and swEF1. Therefore, the construction works for all these notions without any modification.
tEF1.   The only modification for the case of tEF1, we duplicate $G_{1}$ and $G_{2}$. That is, there are two copies of special items specific for each agent.
EFL.   For EFL, the reduction is identical. Observe that each agent values as $t$ the large item he wants and that is allocated to the other agent due to social impact. Hence he has to obtain at least value of $t$ from standard items. ∎
@@
However, the previous hardness result is highly dependent on the fact that the agents’ valuations for the items are exponential in the input size. That is, the result shows only weak NP-hardness. In the following, we complement the hardness and show that if the number of agents is a fixed constant and the valuations are polynomially bounded in the input size, then there is an efficient algorithm. Notably, the algorithm is not based on an explicit DP formulation, as one might expect, but rather models our problem as graph reachability. Thanks to this, the algorithm can be implemented very efficiently by generating the graph on-demand and finding the solution using, e.g., a variant of the A algorithm (HartNR1968, 33, 24, 39, 51).
@@
Proof.
Let $\nu=\max_{i\in{{{N}}}}\max_{g\in{M}}|v_{i}(g)|$} be the largest value an agent can receive from a single item. We build a directed acyclic graph $G=(V,E)$ with m+2{m}+2 layers of vertices $V_{s}=\{s\},V_{t}=\{t\},\{V^{i}\mid i\in{M}\}$. The vertex $s$ is the source and $t$ is the sink of the graph. Vertices in the layer $V^{i}$ are in one-to-one correspondence with the set
$\{0,\ldots,\nu\cdot{m}\}^{{{N}}\times{{N}}}\times\{0,\ldots,\nu\}^{{{N}}\times{{N}}}\,$,
that is, the product of
  1. (1)
    a set of ${{n}}^{2}$-sized tuples of numbers in $\nu\cdot{m}]_{0}$ (this way, we store the value of the bundles as seen by all the agents), and
  2. (2)
    a set of ${{n}}^{2}$-sized tuples of numbers in $\nu]_{0}$ (we store the value of a largest item as seen by all the agents).
The intended meaning of such a vertex $v^{i}_{\bm{x},\bm{y}}$ is that we are having a partial allocation assigning items in M{M} preceding $i$ in some fixed order in such a way that the value $x_{a,b}={v}_{a}({A}_{b})$ (note that we allow $a=b$) and the value of a best item in the bundle ${A}_{b}$ of $b\in{{N}}$ as seen by $a\in{{N}}$ is $y_{a,b}$. We interpret the source vertex $s$ as the vertex $v^{0}_{\bm{0},\bm{0}}$, i.e., no items assigned and all the values set to 0.
We begin the exposition for the objective EF1. We construct the arcs of the graph in a layer by layer fashion. The arcs are always between consecutive layers. Let $i\in{M}$ be an item, we construct the arcs from the previous layer $V^{j}$ to the layer $V^{i}$ as follows. For each vertex $v^{j}_{\bm{x},\bm{y}}$ we add at most n{{n}} arcs corresponding to assigning the item $i$ to each agent. That is, for each agent $c\in{{N}}$ we add the arc $(v^{j}_{\bm{x}},v^{j}_{\bm{\widehat{x}},\bm{\widehat{y}}})$ if $c$ is an agent maximizing the social impact if $i$ is assigned to them,
$\widehat{x}_{a,b}=\begin{cases}x_{a,b}&a,b\in{{N}}\setminus\{c\}\\ x_{c,a}&a\in{{N}}\setminus\{c\}\\ x_{a,c}+{v}_{a}(i)&a\in{{N}}\end{cases}\,$,
in other words, all the agents add their value of $i$ to how they perceive the bundle ${A}_{c}$ of agent $c$ (including $c$); and we update the value $y_{a,b}$ based on if $i$ is a best item in the bundle of $c$ as viewed by the other agents
$\widehat{y}_{a}=\begin{cases}y_{a,b}&a\in{{N}},b\in{{N}}\setminus\{c\}\\ \max\{y_{a,c},{v}_{a}(i)\}&a\in{{N}}\end{cases}\,.$
The arcs from the last layer to the sink $t$ we add from all vertices corresponding to EF1 allocations, that is, if for all pairs of agents $a,b\in{{N}}$ the following holds for the vertex $v_{\bm{x},\bm{y}}$ in the last layer:
$x_{a,a}\geq x_{a,b}-y_{a,b}\,$,
in other words, if we remove one of the most valuable items (with the value $y_{a,b}$) from ${A}_{b}$ (which is valued by $a$ as $x_{a,b}$), then $a$ does not envy to $b$ (since $a$ values their bundle to $x_{a,a}$). Now, it follows from the construction that any $s$-$t$-path in $G$ corresponds to allocation of all items to agents that maximizes the social impact and is EF1. Furthermore, $G$ is acyclic, since we always increase the number of the layer (and never go back). To finish the argument, we discuss the size of $G$. The number of vertices of $G$ is
$2+$ $|[\nu\cdot{m}]_{0}^{{{N}}\times{{N}}}\times[\nu]_{0}^{{{N}}\times{{N}}}|\cdot{m}$
$=2+\left((1+|\nu|\cdot{m})^{{{n}}^{2}}+((1+\nu)^{{{n}}^{2}})\right)\cdot{m}$
a size that is pseudopolynomial if n{{n}} is a constant. Note that we can find an $s$-$t$-path if it exists using, e.g., BFS in linear-time in the size of the graph. We stress here that even though the size of the graph $G$ is clearly impractical even for mild settings of $\nu,{{n}},{m}$, we can generate it in an on-demand fashion and find the solution quickly in practice using, e.g., the A algorithm.
wEF1. If we change the objective from EF1 to wEF1, the only affected part of the construction of the graph $G$ are the edges towards the sink $t$. This is the only place which is objective-dependent. The arcs from the last layer to the sink $t$ we add from all vertices corresponding to wEF1 allocations, that is, if for all pairs of agents $a,b\in{{N}}$ the following holds for the vertex $v_{\bm{x},\bm{y}}$ in the last layer:
$\frac{x_{a,a}}{w_{a}}\geq\frac{x_{a,b}-y_{a,b}}{w_{b}}\,$,
[…]
$\widehat{y}_{a}=\begin{cases}y_{a,b}&a\in{{N}},b\in{{N}}\setminus\{c\}\\ {v}_{a}(i)&a\in{{N}}\end{cases}\,.$
The rest is done in the same way as in EF1.
tEF1. We verify that
$x_{a,a}+y_{a,b}\geq x_{a,b}-y_{a,b}\,$,
when adding arcs to $t$, since this corresponds not only to $b$ not having the item of $a$-value $y_{a,b}$ but at the same time increasing the value of $a$’s bundle by adding this item to their bundle.
EF. In this case, we can work with $V^{i}_{\bm{x}}$ only, since we do not need to remove any items and therefore it suffices to track the values of the bundles.
EFL. This case is similar to the sEF1 case, that is, we add two arcs corresponding to updating the vector $\bm{y}$ or not. Finally, when adding the arcs to the sink $t$, we have to check
$x_{a,a}\geq x_{a,b}-y_{a,b}\quad\text{and}\quad y_{a,b}\leq x_{a,a}\,$,
$x_{a,a}\geq x_{a,b}\,$
or
$x_{a,b}=y_{a,b}$
holds. ∎
4. Socially Aware Agents
@@
Motivated by the experimental work of (HerreinerP2009, 34) and (HosseiniKSVX2025, 36), which explore what real-life users perceive to be fair outcomes, (Hosseini2024, 35) argues that “developing fair algorithms may require axioms that are able to capture solutions that take the social context into account beyond perceived envy.” In this section, we build up on this idea and follow the approach introduced by (FlamminiGV2025, 29) and assume socially aware agents. Intuitively, an agent is socially aware if they are willing to accept a formally unfair outcome in case it leads to a greater social impact.
Formally, we define envy-freeness with socially aware agents as follows.
Definition 0.
We say that an allocation $\mathcal{A}$} is envy-free with socially aware agents (SA-EF), if for each pair of agents $i,j\in{{N}}$, at least one of the following conditions holds:
${v}_{i}({A}_{i})\geq{v}_{i}({A}_{j})$ or $s_{i}({A}_{j})<s_{j}({A}_{j})\,.$
Based on the above definition, we will say that an agent $i$ SA-envies $j$, if and only if ${v}_{i}({A}_{i})<{v}_{i}({A}_{j})\text{ and }s_{i}({A}_{j})\geq s_{j}({A}_{j})$. Notice that if the social impact of all agents is the same for all agents, SA-EF simply reduces to EF, a computationally notoriously hard fairness notion which is rarely satisfiable. Henceforth, in the rest of the section, we focus on natural relaxations of EF.
@@
Proof.
In any SIM allocation, for each pair of agents $i$ and $j$ we have ${s}_{i}({A}_{j})\leq{s}_{j}({A}_{j})$, so the second condition of the definition (with $\leq$) is always satisfied. ∎
In their paper, (FlamminiGV2025, 29) proved that there is a modification of the famous envy-cycle elimination algorithm (LiptonMMS2004, 41) that always finds an SA-EF1 allocation. We extend their result in two ways. In our first result, we show that SA-swEF1 allocations (and hence also SA-wEF1 allocations) are also guaranteed to exist. Interestingly, our approach is based on a picking sequence similar to the round-robin algorithm. More precisely, the algorithm is a modification of the weighted picking sequence protocol proposed by (ChakrabortyISZ2021, 21), which reduces to a round-robin approach in the case of identical agent weights. However, we only allow agents to pick goods for which they maximize social impact, and we force agents to skip their turn whenever they do not maximize the social impact of any unallocated good. The proof then basically follows from (ChakrabortyISZ2021, 21) and a simple observation that agent $i$ can SA-envy agent $j$ only if $j$ chose only goods for which $i$ maximizes social impact. Therefore, from the point of view of $i$, we can completely ignore all agents that get an item for which $i$ does not maximize social impact. On the other hand, if we consider an auxiliary instance in which agents only value items for which they maximize impact and get a dummy item that everyone values 0 whenever they skip a turn we get an allocation that is swEF1 according to (ChakrabortyISZ2021, 21) and in which: 1) All agents get same bundle plus some dummy items. 2) Every agent values every bundle for which the agent maximizes the social impact and all non-dummy items in it exactly the same in both instance.